OpenStudy (anonymous):

-(pi/4)sin(pix/4) solve for x, x >(or equal to) 0 x <(or equal to) 10 and i should end up with x= 0,4,8 how does that work?

OpenStudy (dumbcow):

is there any other info? i think it should be equal to something in order to solve for x

OpenStudy (anonymous):

equal to 0 sorry

OpenStudy (dumbcow):

wait nevermind its -(pi/4)sin(pi*x/4) = 0 divide by -pi/4 on both sides leaving sin(pi*x/4) = 0 now think about the graph of the sin function it is only zero at x=0,pi,2pi...

OpenStudy (anonymous):

ohhhh but then i get where it is 0. im not gettting that once i get to sin(pi*x/4) what does the pi/4 contribute? idk if im stating that correctly

OpenStudy (dumbcow):

think of it this way say y = pi*x/4 we know sin(y) = 0 at y=0,pi,2pi so set up three equation pix/4 = 0 or pi or 2pi and solve for x

OpenStudy (anonymous):

oh ok, thats what i did wrong. i solved for 0, but i forgot that sin is 0 at pi and 2pi. and that i need to account for that too. thank you sooo much!

OpenStudy (dumbcow):

your welcome :)